OpenStudy (anonymous):

is the absolute value of 1-1/5x 1+1/5x or does it stay 1-1/5x?

OpenStudy (anonymous):

it is neither

OpenStudy (anonymous):

\[|1-\frac{1}{5}x|\] is positive but that does not make it \[1+\frac{1}{5}x\]

OpenStudy (anonymous):

\[|1-\frac{1}{5}x| = \left\{\begin{array}{rcc} 1-\frac{1}{5}x & \text{if} & x \leq 5 \\ \frac{1}{5}x-1& \text{if} & x>5\\ \end{array} \right.\]